OpenStudy (marmar10):

Help !! solve the following equations for x over the interval [0, 2pi]

OpenStudy (anonymous):

wat

OpenStudy (marmar10):

\[\sin2x + \sqrt{2}sinx=0 \] \[2 \cos ^{2}x+1 = 3cosx\]

OpenStudy (marmar10):

cos2x+2=sinx \[2\sin ^{2}x-sinx=1\] If someone could help out that would be great :) Im a little confused how to start off

OpenStudy (campbell_st):

so looking at th 1st equation youcan use substitution for sin(2x) a sin(2x) = sin(x)cos(x) so the equation becomes \[2\sin(x)\cos(x)+ \sqrt{2}\sin(x) = 0\] so te equation can now be factored then solved.

OpenStudy (campbell_st):

rewrite the 2nd equation as \[2\cos^2(x) - 3\cos(x) +1 = 0\] you now have a quadratic equation that can be solvd by factoring

OpenStudy (campbell_st):

\[2\sin^2(x) - \sin(x) - 1 = 0\] another quadratc equation that can be solved by factoring..

OpenStudy (marmar10):

thank you that helps a lot :) for the first one would i have to to square the whole equation to get rid of the square root right?

OpenStudy (campbell_st):

cos(2x) + 2 = sin(x) use substitution \[\cos(2x) \cos^2(x) - \sin^2(x)\] then writing evething in ters of sin(x) \[\cos(2x) = 1- 2\sin^2(x)\] so the equation can be written as \[2\sin^2x + \sin(x) - 3 =0\] whicis agin a quadratic quation tha can be solved by factoring

OpenStudy (campbell_st):

for the 1st equation remove the common factor of sin(x) \[\sin(x)(2\cos(x) - \sqrt{2}) = 0\]

OpenStudy (marmar10):

ah ok thank you very much :)

OpenStudy (anonymous):

marmar 10 want a medal?

OpenStudy (anonymous):



OpenStudy (anonymous):

times up you get a medal